OpenStudy (anonymous):

In solving for the solution of a non homogeneous differential equation using the method of undetermined coefficient, example, y'' + y' + y = 5. Yp will be guessed as a constant. Is that right?

OpenStudy (turingtest):

that depends on the solution to the complimentary

OpenStudy (turingtest):

but I guess so in this case

OpenStudy (turingtest):

which makes finding it pretty darn easy

OpenStudy (anonymous):

thanks :D well my problem came from the this

OpenStudy (anonymous):

y''' - y'' = e^x + 4

OpenStudy (anonymous):

My answer is y = yh + yp = A + Bx + Ce^x + Dxe^x + E, but my friend told me to omit the E

OpenStudy (turingtest):

yeah, A+E is just another constant, so you can call it A again I think your homogeneous solution is wrong though

OpenStudy (turingtest):

I am confused, is this the same problem?

OpenStudy (anonymous):

my above answer is the answer to the y''' - y'' = e^x + 4

OpenStudy (turingtest):

you mean that y'' + y' + y = 5 is your answer to y''' - y'' = e^x + 4 ?? now I'm more confused

OpenStudy (anonymous):

sorry the y'' + y' + 5 was a prelimary question. It is not related to the stuff in the comments

OpenStudy (turingtest):

ohhhhhhhkay so you meant that y = yh + yp = A + Bx + Ce^x + Dxe^x is your guess to y''' - y'' = e^x + 4 ???

OpenStudy (anonymous):

yea haha sorry for the confusion

OpenStudy (turingtest):

now problem, but I think your guess for the particular is wrong, though I am having trouble finding it myself too

OpenStudy (turingtest):

what did you get for yh?

OpenStudy (anonymous):

A+ Bx+ Ce^x

OpenStudy (turingtest):

ok, that's right (though I would call it yh=c1+c2x+c3e^x to avoid confusion) and what did you guess for yp ?

OpenStudy (anonymous):

Dxe^x + E , but E is a constant so same as A + E

OpenStudy (turingtest):

that's why I said to avoid confusion, A and E are different kinds of constants. A can only be found if given an initial condition, E can be found without one

OpenStudy (anonymous):

hmm so in this case i should not omit the E?

OpenStudy (turingtest):

so call your guess Axe^x+B but that is wrong why? because no part of the particular solution can be linearly dependent on the complimentary since c1 and A are both constants, we have a problem; we need to multiply by x until we get something linearly independet

OpenStudy (turingtest):

in this case, no do not omit the E, we must make it linearly independent

OpenStudy (anonymous):

wow thanks a lot for the explanation :D now i understand :D

OpenStudy (anonymous):

so my answer will be y=C + C1x + C2e^x + Axe^x + Bx^2

OpenStudy (anonymous):

oops its C1 from the start

OpenStudy (turingtest):

almost, but rememeber that I said you can find the values of A and B

OpenStudy (turingtest):

plug your guess for the particular into the differential equation and what do you get?

OpenStudy (anonymous):

i need to differentiate the particular like three times and plug in

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

hmm my question paper said there was no need to determine any coefficients

OpenStudy (turingtest):

well, you certainly *can* find A and B if you want to, but if your paper says you don't have to find any constants at all, so be it.

OpenStudy (anonymous):

yea, i think i could manage the finding of A and B if needed :D. Thanks for your help again :D

OpenStudy (anonymous):

i guess ill be closing this question. Thanks for your help!

OpenStudy (turingtest):

welcome :)